A balloon contains 46.3 L of Helium at 42 oC, and 2.01 atm is released into the air. At a certain altitude, the temperature falls to 21 oC, and the pressure falls to 0.25 atm. What is the volume of the balloon under these conditions?


A. 3.5 L

B. 35.0 L

C. 350 L

D. 3500 L

E. None of the Above


1
Expert's answer
2021-05-18T07:52:25-0400

We have two conditions:


First condition:

"V_1=46.3l"

"T_1=42^0 C"

"P1=2.01atm"


Second condition:

"letV_2=V_2"

"T_2=21^0C"

"P_2=0.25atm"


Now by the help of gas equation:


"\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"


"\dfrac{2.01\times46.3}{42}=\dfrac{0.25\times V_2}{21}"


"V_2=186.126l"


So our answer is Option E means none of the above is correct.